# references.bib

### v1

@@ 6 Representations of $S_{3}$ in $Q_{m}(3,\mathbf{F}_{3})$
Suppose that for all $i\leq m$, $Q_{i}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators in degree $d$ and $6i+3-d$ respectively for some $d$. If $K,L$ are distinct generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then there are two other homogeneous generators $K_{1},L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ in the same degrees as $K,L$, respectively such that as a representation of $S_{3}$, $K_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $K$ and $L_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $L$. Moreover, there are no relations between $K_{1},L_{1}$ over the symmetric polynomials, and there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$ that generate a copy of $\mathrm{triv-sign-triv}$.
@@
For the inductive step, let $K^{\prime},L^{\prime}$ be the generators of $Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ and let $K^{\prime}_{1},L^{\prime}_{1}$ be the corresponding generators of $Q_{m-1}(3,\mathbf{F}_{3})$. Without loss of generality, we can choose $K^{\prime}_{1},L^{\prime}_{1}$ to be $s_{23}$-invariant with $(1-s_{12})K^{\prime}_{i}=K^{\prime},(1-s_{12})L^{\prime}_{i}=L^{\prime}$ (similar to in the base case). Let $K,L$ be generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$. Then since $K,L\in Q_{m-1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, we can write $K=P_{1}K^{\prime}+Q_{1}L^{\prime}$, $L=P_{2}K^{\prime}+Q_{2}L^{\prime}$ for symmetric polynomials $P_{1},P_{2},Q_{1},Q_{2}$. Then it follows that $K_{1}:=P_{1}K^{\prime}_{1}+Q_{1}L^{\prime}_{1},L_{1}:=P_{2}K^{\prime}_{1}+Q_{2}L^{\prime}_{1}$ each generate a copy of $\mathrm{triv-sign-triv}$ that contains $K,L$, respectively. Moreover, if there is some relation $P_{3}K_{1}+Q_{3}L_{1}=0$ for symmetric polynomials $P_{3},Q_{3}$, then applying $1-s_{12}$ to this equation would yield $P_{3}K+Q_{3}L=0$, which violates Lemma 4.4.
Next, we show that $K_{1},L_{1}$ are $m$-quasi-invariant. As the computations are the same for both polynomials, we give the proof only for $K_{1}$. First, note that $(1-s_{23})K_{1}=0$ since both $K^{\prime}_{1},L^{\prime}_{1}$ are $s_{23}$-invariant. Next, note that $(1-s_{12})K_{1}=K$ is divisible by $(x_{1}-x_{2})^{2m+1}$ by Lemma 4.1. Finally, note that since $K_{1}$ is $s_{23}$-invariant, we have
@@
Note that $K_{1},L_{1}$ are the minimal degree polynomials such that $(1-s_{12})K_{1},(1-s_{12})L_{1}$ are symmetric polynomial multiples of $K,L$, respectively, so they cannot be generated by any other generators and thus must be a generators themselves. Then assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$ that generates a copy of $\mathrm{triv-sign-triv}$. Then $(1-s_{12})T$ is contained in a copy of $\mathrm{sign-triv}$ and is $s_{12}$-antiinvariant, so we can write $(1-s_{12})T=S_{1}K+S_{2}L$ for symmetric polynomials $S_{1},S_{2}$. Then $T,S_{1}K_{1}+S_{2}L_{1}$ generate copies of $\mathrm{triv-sign-triv}$ with the same $\mathrm{sign-triv}$ submodule, so they generate a copy of
@@
The generators $1,K,K_{1},L,L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ defined in Proposition 2.9, Theorem 5.3, and Proposition 6.1 have no relations between them over the symmetric polynomials.
@@
In the non-modular case, one has that the polynomial $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ is a generator of $Q_{m}(n,\mathbbm{k})$, as it is the lowest degree quasi-invariant in the sign module. However, from Lemma 4.4 we have that in characteristic 3,
$(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i<j}(x_{i}-x_{j})^{2m+1},$
so $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ is not a generator. We can take this calculation further, and note that $(L+s_{23}L)K_{1}-(K+s_{23}K)L_{2}$ would then generate a copy of $\mathrm{triv-sign}$, as the quotient of this module by the space generated by $(L+s_{23}L)K-(K+s_{23}K)L$ must be a trivial module.
Now, we now only need to consider the modules $\mathrm{triv-sign},\mathrm{sign-triv-sign}$. To motivate the results that follow, we start by considering 0-quasi-invariants.
@@
Note that from Corollary 6.2 we know that $Q_{0}(3,\mathbf{F}_{3})$ has 5 generators $1,x_{1}-x_{2},(x_{1}-x_{2})x_{3},x_{1},x_{1}(x_{2}+x_{3})$ with no relations between them. By examining the dimension of the space of all homogeneous degree 3 polynomials, we have that $Q_{0}(3,\mathbf{F}_{3})[3]$ is 10-dimensional, and so far we have accounted for $3+2+2+1+1=9$ dimensions. Moreover, every irreducible representation is accounted for, so this extra dimension must be an extension of an existing indecomposable representation. The only indecomposable representations that have nontrivial extensions are the triv generated by $x_{1}x_{2}x_{3}$ and the $\mathrm{triv-sign}$ generated by
$E:=(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})x_{1}+(x_{1}+x_{2}+x_{3})(x_{1}(x_{2}+x_{3}))=-x_{1}^{2}x_{2}-x_{1}^{2}x_{3}+x_{1}x_{2}^{2}+x_{1}x_{3}^{2}.$
@@
for symmetric polynomials $P_{i}$. Apply $1-s_{12}$ to this equation to get
@@
Finally, note that as $E$ can be written in terms of symmetric polynomial multiples of $x_{1},(x_{2}+x_{3})x_{1}$, this equation would be a relation between the first 5 generators of $Q_{0}(3,\mathbf{F}_{3})$. We have seen this is impossible, so we have $P_{6}=0$, and hence all of the $P_{i}$ must be 0.
Finally, note that the Hilbert series of the submodule of $Q_{0}(3,\mathbf{F}_{3})$ generated by these 6 polynomials is
$\frac{1+2t+2t^{2}+t^{3}}{(1-t)(1-t^{2})(1-t^{3})}=\frac{1}{(1-t)^{3}}.$
This is exactly the Hilbert series of $Q_{0}(3,\mathbf{F}_{3})$, so there are no more generators of $Q_{0}(3,\mathbf{F}_{3})$. ∎
@@
  1. 1)
    Let $T\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{triv-sign}$. Then $T$ is the sum of a symmetric polynomial multiple of $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ and a symmetric polynomial. Conversely, any symmetric polynomial multiple of $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ generates a copy of $\mathrm{triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
  2. 2)
    Let $T_{1}\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{sign-triv-sign}$. Then $T_{1}$ is the sum of a symmetric polynomial multiple of $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ and a symmetric polynomial multiple of $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$. Conversely, any symmetric polynomial multiple of $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ generates a copy of $\mathrm{sign-triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
@@
1) We first prove the lemma for $m=0$. Consider some $T$ as above, and note that $(1-s_{12})T$ is contained in the sign representation, so by Proposition 2.9 we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ for some symmetric polynomial $P$. Then note that $PE,T$ generate two copies of $\mathrm{triv-sign}$ with the same sign subrepresentation, so they generate a copy of
@@
Now, consider general $m$. By the above we have that any $T$ must be of the form $T=PE+Q$ for symmetric polynomials $P,Q$. Then since $T$ is $m$-quasi-invariant, we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ is divisible by $(x_{1}-x_{2})^{2m+1}$. So $P$ is divisible by $(x_{1}-x_{2})^{2m}$, and it must also be divisible by $\prod_{i<j}(x_{i}-x_{j})^{2m}$ since it is symmetric.
@@
2) This proof is similar to part 1). For $m=0$, any $T_{1}$ must have that $(1+s_{23})T_{1}$ is in a $\mathrm{triv-sign}$ representation, so $(1+s_{23})T_{1}=PE$ for some $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$. Then $T_{1},PF$ generate a copy of
@@
which implies the result for $m=0$. Then the extension to general $m$ is the same as in part 1). The converse is clear, as before. ∎
Finally, we can prove Theorem 1.3 for $p=3$. Note that this also implies Conjecture 5.2.
@@
$Q_{m}(3,\mathbf{F}_{3})$ is freely generated by $1$, the two generators $K,L$ of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ from Theorem 5.3, the two generators $K_{1},L_{1}$ from Proposition 6.1, and the generator $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ from Lemma 6.6.
@@
Let us first show that there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$. Assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$. Then $T$ cannot generate a copy of $\mathrm{triv}$ by Proposition 2.9 and it cannot generate a copy of $\mathrm{sign-triv}$ or $\mathrm{triv-sign-triv}$ by Theorem 5.3 and Proposition 6.1. If it generates a copy of $\mathrm{sign}$, then by Proposition 2.9 it must be $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$, but this polynomial is generated by $K,L$ by Lemma 4.4, so it cannot be a generator. If it generates a copy of $\mathrm{triv-sign}$, then it is $E\prod_{i<j}(x_{i}-x_{j})^{2m}$ by Lemma 6.6. But this is generated by $K_{1},L_{1}$ by Remark 6.3. Finally, by Lemma 6.6 the only generator that generates a copy of $\mathrm{sign-triv-sign}$ is $F\prod_{i<j}(x_{i}-x_{j})^{2m}$.
Finally, we show there are no relations between the 6 generators. Note that this also implies $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ is a generator, since it is not generated by the other 5 generators. But this is clear: we already know there are no relations between the first 5 generators by Corollary 6.2. If there was a relation involving $F\prod_{i<j}(x_{i}-x_{j})^{2m}$, then note that since every generator is generated by the generators of $Q_{0}(3,\mathbf{F}_{3})$, this would induce a relation on those generators. Moreover, $F\prod_{i<j}(x_{i}-x_{j})^{2m}$ is the only generator not generated by the first 5 generators of $Q_{0}$, so the induced relation would be nontrivial. But there is no such relation by Proposition 6.5. ∎

### v2

@@ 6 Representations of $\boldsymbol{S_{3}}$ in $\boldsymbol{Q_{m}(3,\mathbf{F}_{3})}$
Suppose that for all $j\leq m$, $Q_{j}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has generators in degree $d$ and $6j+3-d$ respectively for some $d$. If $K$, $L$ are distinct generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then there are two other homogeneous generators $K_{1}$, $L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ in the same degrees as $K$$L$, respectively such that as a representation of $S_{3}$, $K_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $K$ and $L_{1}$ generates a copy of $\mathrm{triv-sign-triv}$ containing $L$. Moreover, there are no relations between $K_{1}$, $L_{1}$ over the symmetric polynomials, and there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$ that generate a copy of $\mathrm{triv-sign-triv}$.
@@
[…]
Next, we show that $K_{1}$, $L_{1}$ are $m$-quasi-invariants. As the computations are the same for both polynomials, we give the proof only for $K_{1}$. First, note that $(1-s_{23})K_{1}=0$ since both $K^{\prime}_{1}$, $L^{\prime}_{1}$ are $s_{23}$-invariant. Next, note that $(1-s_{12})K_{1}=K$ is divisible by $(x_{1}-x_{2})^{2m+1}$ by Lemma 4.1. Finally, note that since $K_{1}$ is $s_{23}$-invariant, we have
@@
Note that $K_{1}$, $L_{1}$ are the minimal degree polynomials such that $(1-s_{12})K_{1}$, $(1-s_{12})L_{1}$ are symmetric polynomial multiples of $K$, $L$, respectively, so they cannot be generated by any other generators and thus must be generators themselves. Then assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$ that generates a copy of $\mathrm{triv-sign-triv}$. Then $(1-s_{12})T$ is contained in a copy of $\mathrm{sign-triv}$ and is $s_{12}$-antiinvariant, so we can write $(1-s_{12})T=S_{1}K+S_{2}L$ for symmetric polynomials $S_{1}$, $S_{2}$. Then $T$, $S_{1}K_{1}+S_{2}L_{1}$ generate copies of $\mathrm{triv-sign-triv}$ with the same $\mathrm{sign-triv}$ submodule, so they generate a copy of
@@
The generators $1$, $K$, $K_{1}$, $L$, $L_{1}$ of $Q_{m}(3,\mathbf{F}_{3})$ defined in Proposition 2.11, Theorem 5.3 and Proposition 6.1 have no relations between them over the symmetric polynomials.
@@
In the non-modular case, one has that the polynomial $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$ is a generator of $Q_{m}(n,\mathbbm{k})$, as it is the lowest degree quasi-invariant in the sign module. However, from Lemma 4.4 we have that in characteristic 3,
$(L+s_{23}L)K-(K+s_{23}K)L=c\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1},$
so $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$ is not a generator. We can take this calculation further, and note that $(L+s_{23}L)K_{1}-(K+s_{23}K)L_{1}$ would then generate a copy of $\mathrm{triv-sign}$, as the quotient of this module by the space generated by $(L+s_{23}L)K-(K+s_{23}K)L$ must be a trivial module.
It remains to consider the modules $\mathrm{triv-sign}$, $\mathrm{sign-triv-sign}$. To motivate the results that follow, we start by considering 0-quasi-invariants.
@@
Note that from Corollary 6.2 we know that $Q_{0}(3,\mathbf{F}_{3})$ has 5 generators $1$, $x_{1}-x_{2}$, $(x_{1}-x_{2})x_{3}$, $x_{1}$, $x_{1}(x_{2}+x_{3})$ with no relations between them. By examining the dimension of the space of all homogeneous degree 3 polynomials, we have that $Q_{0}(3,\mathbf{F}_{3})[3]$ is 10-dimensional. Since $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ is 3-dimensional in degree $3$, 2-dimensional in degree $2$, and 1-dimensional in degree $1$, so far we have accounted for only $3+2+2+1+1=9$ dimensions. Moreover, every irreducible representation is accounted for, so this extra dimension must be an extension of an existing indecomposable representation. The only indecomposable representations that have nontrivial extensions are the $\mathrm{triv}$ generated by $x_{1}x_{2}x_{3}$ and the $\mathrm{triv-sign}$ generated by
$E:={}$ $(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})x_{1}+(x_{1}+x_{2}+x_{3})(x_{1}(x_{2}+x_{3}))$
[…]
@@
for symmetric polynomials $P_{j}$. Apply $1-s_{12}$ to this equation to get
@@
Finally, note that as $E$ can be written in terms of symmetric polynomial multiples of $x_{1}$, ${(x_{2}+x_{3})x_{1}}$, this equation would be a relation between the first 5 generators of $Q_{0}(3,\mathbf{F}_{3})$. We have seen this is impossible, so we have $P_{6}=0$, and hence all of the $P_{j}$ must be 0.
Let $Q_{0}^{\prime}$ be the submodule of $Q_{0}(3,\mathbf{F}_{3})$ generated by these 6 polynomials. Then as the polynomials freely generate $Q_{0}^{\prime}$ as a $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, we have that the Hilbert series of $Q_{0}^{\prime}$ is
$\mathcal{H}(Q_{0}^{\prime})=\big{(}1+2t+2t^{2}+t^{3}\big{)}\mathcal{H}\big{(}\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}\big{)}=\frac{1+2t+2t^{2}+t^{3}}{(1-t)(1-t^{2})(1-t^{3})}=\frac{1}{(1-t)^{3}}$
by the fundamental theorem of symmetric polynomials. This is exactly the Hilbert series of $Q_{0}(3,\mathbf{F}_{3})$, so $Q_{0}^{\prime}=Q_{0}(3,\mathbf{F}_{3})$ and there are no more generators of $Q_{0}(3,\mathbf{F}_{3})$. ∎
@@
  1. $1$
    Let $T\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{triv-sign}$. Then $T$ is the sum of a symmetric polynomial multiple of $E\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ and a symmetric polynomial. Conversely, any symmetric polynomial multiple of $E\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ generates a copy of $\mathrm{triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
  2. $2$
    Let $T_{1}\in Q_{m}(3,\mathbf{F}_{3})$ generate a copy of $\mathrm{sign-triv-sign}$. Then $T_{1}$ is the sum of a symmetric polynomial multiple of $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ and a symmetric polynomial multiple of $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$. Conversely, any symmetric polynomial multiple of $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ generates a copy of $\mathrm{sign-triv-sign}$ in $Q_{m}(3,\mathbf{F}_{3})$.
@@ 2. $2$
1. We first prove the lemma for $m=0$. Consider some $T$ as above, and note that $(1-s_{12})T$ is contained in the sign representation, so by Proposition 2.11 we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ for some symmetric polynomial $P$. Then note that $PE$, $T$ generate two copies of $\mathrm{triv-sign}$ with the same sign subrepresentation, so they generate a copy of
@@
Now, consider general $m$. By the above we have that any $T$ must be of the form $T=PE+Q$ for symmetric polynomials $P$, $Q$. Then since $T$ is $m$-quasi-invariant, we have $(1-s_{12})T=P(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3})$ is divisible by $(x_{1}-x_{2})^{2m+1}$. So $P$ is divisible by $(x_{1}-x_{2})^{2m}$, and it must also be divisible by $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ since it is symmetric.
@@
[…]
@@
which implies the result for $m=0$. Then the extension to general $m$ is the same as in part (1). The converse is clear, as before. ∎
Finally, we can prove Theorem 1.3 for $p=3$.
@@
$Q_{m}(3,\mathbf{F}_{3})$ is freely generated by $1$, the two generators $K$, $L$ from Theorem 5.3, the two generators $K_{1}$, $L_{1}$ from Proposition 6.1, and the generator $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ from Lemma 6.6.
@@
Let us first show that there are no other generators of $Q_{m}(3,\mathbf{F}_{3})$. Assume for contradiction that there is some other generator $T$ of $Q_{m}(3,\mathbf{F}_{3})$. Then $T$ cannot generate a copy of $\mathrm{triv}$ by Proposition 2.11 and it cannot generate a copy of $\mathrm{sign-triv}$ or $\mathrm{triv-sign-triv}$ by Theorem 5.3 and Proposition 6.1. If it generates a copy of $\mathrm{sign}$, then by Proposition 2.11 it must be $\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m+1}$, but this polynomial is generated by $K$, $L$ by Lemma 4.4, so it cannot be a generator. If it generates a copy of $\mathrm{triv-sign}$, then it is $E\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ by Lemma 6.6. But this is generated by $K_{1}$, $L_{1}$ by Remark 6.3. Finally, by Lemma 6.6 the only generator that generates a copy of $\mathrm{sign-triv-sign}$ is $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$.
Finally, we show there are no relations between the 6 generators. Note that this also implies $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ is a generator, since it is not generated by the other 5 generators. But this is clear: we already know there are no relations between the first 5 generators by Corollary 6.2. If there was a relation involving $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$, then note that since every generator is generated by the generators of $Q_{0}(3,\mathbf{F}_{3})=\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$, this would induce a relation on those generators. Moreover, the generators other than $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ each generate a copy of an indecomposable representation that is not $\mathrm{sign-triv-sign}$, so they are each generated by the first 5 generators of $Q_{0}(3,\mathbf{F}_{3})$. Meanwhile, $F\prod_{i_{1}<i_{2}}(x_{i_{1}}-x_{i_{2}})^{2m}$ is the only generator not generated by the first 5 generators, so the induced relation would be nontrivial. But there is no such relation by Proposition 6.5. ∎
Note that these generators imply a Hilbert series that agrees with Theorem 1.3 since $K$ is either a minimal degree Ren–Xu counterexample or has degree $13m+$ 1 if one does not exist. In this way, the Hilbert series of $Q_{m}(3,\mathbf{F}_{3})$ agrees with that of $Q_{m}(3,\mathbf{Q})$ if and only if there does not exist a Ren–Xu counterexample. Ren–Xu counterexamples only exist when the conditions of Conjecture 5.2 are satisfied, so Conjecture 5.2 is also implied.